# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 4       $k\leftarrow k+1$;
The rest of the proof follows identically to Lemma 2.6. Then we can use this generalized Lemma to obtain Lemma 2.7 without any additional work, since other than using the bounds in Lemma 2.6, normality is not required. ∎
Another property we will need is the monotonicity of the local metric entropy.
Lemma 2.9.
The function $\varepsilon\mapsto M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)$ is monotone non-increasing.
{toappendix}
Proof of Lemma 2.9.
We mimic the procedure in (lecamshamindramatey2022, Lemma 9). First, observe that it suffices to prove for any $h\in\mathcal{F}$ that $\varepsilon\mapsto M(\varepsilon/c,B(h,\varepsilon)\cap\mathcal{F})$ is monotone non-increasing. To see this, suppose $\varepsilon<\varepsilon^{\prime}$ and that we have proven the monotonicity for each fixed $h\in\mathcal{F}$. Suppose to the contrary that $M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon)<M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon^{\prime}).$ Then there is some $\widetilde{h}\in\mathcal{F}$ such that
$M(\varepsilon^{\prime}/c,B(\widetilde{h},\varepsilon^{\prime})\cap\mathcal{F})>M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon)=\sup_{h\in\mathcal{F}}M(\varepsilon/c,B(h,\varepsilon)\cap\mathcal{F})\geq M(\varepsilon/c,B(\widetilde{h},\varepsilon)\cap\mathcal{F}),$
violating our monotonicity property with $\widetilde{h}$.
Now fix $h\in\mathcal{F}$ and let us prove $\varepsilon\mapsto M(\varepsilon/c,B(h,\varepsilon)\cap\mathcal{F})$ is monotone non-increasing. Suppose $\varepsilon<\varepsilon^{\prime}$. Let $M=M(\varepsilon^{\prime}/c,B(h,\varepsilon^{\prime})\cap\mathcal{F})$ and let $f_{1},\dots,f_{M}$ be a maximal $\varepsilon^{\prime}/c$-packing of $B(h,\varepsilon^{\prime})\cap\mathcal{F}$. We must now construct a $\varepsilon/c$-packing of cardinality at least $M$ of the set $B(h,\varepsilon)\cap\mathcal{F}$. Our strategy will be to pull the chosen packing functions closer to $h$. Namely, we set $g_{i}=\left(1-\frac{\varepsilon}{\varepsilon^{\prime}}\right)h+\frac{\varepsilon}{\varepsilon^{\prime}}f_{i},$ noting that $g_{i}\in\mathcal{F}$ for all $i\in[M]$ by convexity of $\mathcal{F}$. Further, each $g_{i}\in B(h,\varepsilon)$ since
$\|g_{i}-h\|_{L_{2}(\mathbb{P}_{X})}$ $=\left\|\frac{\varepsilon}{\varepsilon^{\prime}}\cdot f_{i}-\frac{\varepsilon}{\varepsilon^{\prime}}\cdot h\right\|_{L_{2}(\mathbb{P}_{X})}=\frac{\varepsilon}{\varepsilon^{\prime}}\cdot\|f_{i}-h\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{\varepsilon}{\varepsilon^{\prime}}\cdot\varepsilon^{\prime}=\varepsilon,$
where we used the fact that the $f_{i}\in B(h,\varepsilon^{\prime})$. Lastly, for $i\neq j$, since the $f_{i}$ are $\varepsilon^{\prime}/c$-separated,
$\|g_{i}-g_{j}\|_{L_{2}(\mathbb{P}_{X})}$ $=\left\|\frac{\varepsilon}{\varepsilon^{\prime}}(f_{i}-f_{j})\right\|_{L_{2}(\mathbb{P}_{X})}>\frac{\varepsilon}{\varepsilon^{\prime}}\cdot\frac{\varepsilon^{\prime}}{c}=\frac{\varepsilon}{c}.$
Thus $g_{1},\dots,g_{M}$ is indeed a $\varepsilon/c$-packing set of $B(h,\varepsilon)\cap\mathcal{F}$ of cardinality $M$. Hence the maximal such packing set has cardinality at least $M$. ∎
{toappendix}
The following technical lemma will be useful for the proof of Theorem 2.11.
Lemma 2.10.
For any integer $J>1$ and constant $\rho>0$, we have
$\sum_{j=2}^{J}\exp\left(-\rho\cdot\frac{1}{2^{2(j-1)}}\right)\leq\frac{\exp\left(-\rho\cdot\frac{1}{2^{2(J-1)}}\cdot\right)}{1-\exp\left(-\rho\cdot\frac{1}{2^{2(J-1)}}\cdot\right)}.$
Proof of Lemma 2.10.
Note first that $\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\in(0,1)$. Then
$\sum_{j=2}^{J}\exp\left(-\rho\cdot\frac{1}{2^{2(j-1)}}\right)$ $=\sum_{j=1}^{J-1}\exp\left(-\rho\cdot\frac{1}{4^{j}}\right)$
$=\sum_{j=1}^{J-1}\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\frac{1}{4^{j+1-J}}\right)$
$=\sum_{j=1}^{J-1}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{\frac{1}{4^{j+1-J}}}$
$=\sum_{j=1}^{J-1}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{4^{J-1-j}}$
$=\sum_{k=0}^{J-2}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{4^{k}}$
$\leq\sum_{k=0}^{\infty}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)\right]^{4^{k}}$
$\leq\sum_{k=1}^{\infty}\left[\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\right)\right]^{k}$
$=\frac{\exp\left(-\rho\cdot\frac{1}{4^{J-1}}\cdot\right)}{1-\exp\left(\rho\cdot\frac{1}{4^{J-1}}\right)}.$
Now we return to our algorithm and quantify its error in the $L_{2}(\mathbb{P}_{X})$-distance. We then show in the subsequent theorem that the minimum of this error and the diameter is in fact the minimax rate. By Remark 2.8, the proof holds for sub-Gaussian noise as well.
Theorem 2.11.
Let $J^{\ast}$ be the maximal integer $J\geq 1$ with $\varepsilon_{J}=\frac{d\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}}{2^{J-2}c}$ satisfying
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{\varepsilon_{J}\cdot c}{\sqrt{L(c/2-1,\sigma,B_{\mathcal{F}})}},c\right)\vee\log 2,$ (2.6)
and $J^{\ast}=1$ if this condition is never satisfied. Let $f^{\ast}(\vec{X},\vec{Y})$ be the output of $J^{\ast}$ steps of Algorithm 1. Then $\mathbb{E}_{X,Y}\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq\overline{C}{\varepsilon_{J^{\ast}}}^{2}$ for some constant $\overline{C}$ depending only on universal constants, $B_{\mathcal{F}}$, and $\sigma$. We defined $L(C,\sigma,B_{\mathcal{F}})$ in (2.3), where $C$ is some sufficiently large universal constant and $c:=2(C+1)$ is the constant we use in the definition of local metric entropy.
{toappendix}
@@
Suppose $J>1$ steps of the algorithm are performed and consider $2\leq j\leq J$. Let $\Upsilon_{j+1}$ be the function added to $\Upsilon$ in step $j$ (e.g., in the first iteration of the loop, we append $\Upsilon_{2}$). During step $j-1$, we start with some $\Upsilon_{j-1}$ and then form a $\frac{d}{2^{j-1}(C+1)}$-packing $M_{j-1}$ of the set $B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$. We then set $\Upsilon_{j}=\mathop{\mathrm{argmin}}_{f\in M_{j-1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$. Applying Lemma 2.7 with $\delta=\frac{d}{2^{j-1}(C+1)}$, $\mathcal{F^{\prime}}:=B(\Upsilon_{j-1},d/2^{j-2})\cap\mathcal{F}$, and conditioning on $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq d/2^{j-2}$ (which implies $\bar{f}\in\mathcal{F}^{\prime}$), we obtain
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}},\Upsilon_{j-1}\right)\leq 3\cdot|M_{j-1}|\exp(-n\varepsilon_{j}^{2}),$
where we define
$\varepsilon_{j}$ $=\frac{d\sqrt{L(C,\sigma,B_{\mathcal{F}})}}{2^{j-1}(C+1)}.$
Further, define $a_{j}=\exp(-n\varepsilon_{j}^{2})$. Observe that $|M_{j-1}|$ is upper-bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{j-2},2(C+1))$, and by Lemma 2.9 this quantity is bounded by $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))$. The probability upper bound does not depend on $\Upsilon_{j-1}$, so we have proven that
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}\bigg{|}\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{j-2}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\exp(-n\varepsilon_{j}^{2}).$
If we set $A_{j}$ to be the event that $\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{j-1}}$, then we can use the property that $P(A_{J})\leq P(A_{J}|A_{J-1}^{c})+P(A_{j-1}|A_{j-2}^{c})+\dots+P(A_{3}|A_{2}^{c})+P(A_{2})$ for each $i$ to conclude that
@@
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\sum_{j=2}^{J}a_{j}$
$\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\frac{a_{J}}{1-a_{J}}.$
The final inequality used Lemma 2.10. Note that $a_{J}<1$ since we assume $C$ is large enough that $L(C,\sigma,B_{\mathcal{F}})>0$.
Note that
$\frac{d}{2^{J-2}}=\frac{2\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}.$
Now, if $n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},2(C+1)\right)$ and $a_{J}=\exp\left(-n\varepsilon_{J}^{2}\right)<\frac{1}{2}$, which implies $(1-a_{J})^{-1}<2$, then we have
@@
$M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}$ $=M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},2(C+1)\right)\cdot\frac{a_{J}}{1-a_{J}}$
$<\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\frac{a_{J}}{1-a_{J}}$
$\leq\exp\left(\frac{n\varepsilon_{J}^{2}}{2}\right)\cdot\exp\left(-n\varepsilon_{J}^{2}\right)\cdot\frac{1}{1-a_{J}}$
$\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Since $a_{J}<\frac{1}{2}$ iff $n\varepsilon_{J}^{2}>\log 2$, we conclude that if
@@
$n\varepsilon_{J}^{2}>2\log M_{\mathcal{F}}^{\operatorname{loc}}\left(\frac{2\cdot\varepsilon_{J}(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}},2(C+1)\right)\vee\log 2,$
@@
then $M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}\leq 2\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right)$.
To summarize, if condition (2.6) holds for $\varepsilon_{J}$, then we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\geq\frac{d}{2^{J-1}}\right)$ $\leq 3\cdot\mathbbm{1}(J>1)M_{\mathcal{F}}^{\operatorname{loc}}(d/2^{J-2},2(C+1))\cdot\frac{a_{J}}{1-a_{J}}$
$\leq 6\cdot\mathbbm{1}(J>1)\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
In other words, (2.6) gives us a lower bound on $P(A_{J}^{c}).$
Now, set $\kappa=\frac{3(C+1)}{\sqrt{L(C,\sigma,B_{\mathcal{F}})}}$. Suppose $J$ satisfies condition (2.6). Then $J\leq J^{\ast}$ with $f^{\ast}$ the output of $J^{\ast}$ steps of the algorithm. By the Cauchy sequence argument in (2.2), we have $\|f^{\ast}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-2}}$. Now if we assume further that $\|\Upsilon_{J}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{d}{2^{J-1}}$, i.e., the event $A_{J}^{c}$, we have by the triangle inequality that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\frac{3d}{2^{J-1}}=\kappa\varepsilon_{J}$. If we call $B_{J}$ the event that $\|f^{\ast}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\kappa\varepsilon_{J}$, we have $P(A_{J}^{c})\leq P(B_{J})$ for $J$ satisfying (2.6). We already obtained a lower bound on $P(A_{J}^{c})$ under this condition, so in summary, for all $J$ satisfying (2.6), we have a lower bound on $P(B_{J})$, namely, that
$\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\kappa\cdot\varepsilon_{J}$
with probability at least
@@
$1-6\cdot\exp\left(-\frac{n\varepsilon_{J}^{2}}{2}\right).$
Let $J^{\ast}$ be the maximal $J$ for which conditions (2.6) all hold, with $J^{\ast}=1$ if none do. Note that $\varepsilon_{J}\downarrow 0$ as $J\to\infty$, so (2.6) will not hold for all sufficiently large $J$ (the left-side of (2.6) is increasing while the right-hand side is non-increasing). This implies $J^{\ast}<\infty$. Further, set $\underline{C}=6$, $C^{\prime}=\frac{1}{2}$. Then for any $1\leq J\leq J^{\ast}$ including $J^{\ast}=1$, we have
$\mathbb{P}_{\bar{f}}\left(\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}\geq\kappa\cdot\varepsilon_{J}\right)$ $\leq\underline{C}\cdot\mathbbm{1}(J>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$
$\leq\underline{C}\cdot\mathbbm{1}(J^{\ast}>1)\exp(-C^{\prime}n\varepsilon_{J}^{2})$ (2.7)
Note the first inequality trivially holds for $J=1$ since $\kappa\varepsilon_{1}=3d$, so that the probability will be 0 as will the right-hand side. Similarly, we extend this to hold for all $J\in\mathbb{Z}$ with $J\leq J^{\ast}$, as if $J\leq 0$, then $\kappa\cdot\varepsilon_{J}=\frac{3d}{2^{J-1}}\geq 6d$. But since $d$ is the diameter of $\mathcal{F}$, the above probability is 00 which is less than than the right hand side of (2.7).

### v5

@@ Appendix B Proofs for Section 3
The rest of the proof follows identically to Lemma 3.8. Then we can use this generalized Lemma to obtain Lemmas 3.9 and 3.10 without any additional work, since other than using the bounds in Lemma 3.8, normality is not required. ∎
@@
If $J^{\ast}=1$, then $\varepsilon_{J^{\ast}}\asymp d$ and the risk bound trivially holds since $d$ is the $L_{2}(\mathbb{P}_{X})$ diameter. Suppose henceforth $J^{\ast}>1$.
Define the event $A_{j}=\{\|\Upsilon_{j}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\tfrac{d}{2^{j-1}}\}$. Observe that $\mathbb{P}(A_{1})=0$ by definition of the diameter, and basic set inclusion properties imply $\mathbb{P}(A_{J})\leq\mathbb{P}(A_{1})+\sum_{j=2}^{J}\mathbb{P}(A_{j}\cap A_{j-1}^{c})$. Let use now bound the probability of each of these intersections.
Part 1: Bounding $\mathbb{P}(A_{j}\cap A_{j-1}^{c})$ for $j\geq 2$.
We start with $j\geq 3$. Set $\delta=\tfrac{d}{2^{j-1}(C+1)}$. We claim the following holds:
$\mathbb{P}(A_{j}\cap A_{j-1}^{c})$
$=\mathbb{P}(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}>\tfrac{d}{2^{j-1}},\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\tfrac{d}{2^{j-2}})$
$\leq\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}\mathbb{P}(\|\bar{f}-\Upsilon_{j}\|_{L_{2}(\mathbb{P}_{X})}>\tfrac{d}{2^{j-1}},\Upsilon_{j-1}=g)$
$=\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}\mathbb{P}(\|\bar{f}-f_{i^{\ast}}\|_{L_{2}(\mathbb{P}_{X})}>(C+1)\delta,\Upsilon_{j-1}=g)$
$\leq\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}\mathbb{P}(\|\bar{f}-f_{i^{\ast}}\|_{L_{2}(\mathbb{P}_{X})}>(C+1)\delta).$ (B.7)
The second line applies a union bound: if $\|\bar{f}-\Upsilon_{j-1}\|_{L_{2}(\mathbb{P}_{X})}\leq\tfrac{d}{2^{j-2}}$, then by definition of $\mathcal{L}(j-1)$ as the $(j-1)$th level of our tree, $\Upsilon_{j-1}=g$ where $g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})$. For the third line, we substitute the definition of $\delta$ and then recall $\Upsilon_{j}=f_{i^{\ast}}$ where $i^{\ast}=\mathop{\mathrm{argmin}}_{i\in[M]}\sum_{j=1}^{n}(Y_{j}-f_{i}(X_{j}))^{2}$ and the $\mathcal{O}(g)=\{f_{1},\dots,f_{M}\}$. We then drop the intersection with $\Upsilon_{j-1}=g$ for the final inequality.
Now, using the definition of $\mathcal{O}(g)$ and Lemma 3.1, note that $f_{1},\dots,f_{M}$ form a $\tfrac{d}{2^{j-2}c}=\tfrac{d}{2^{j-1}(C+1)}=\delta$-covering of the set $\mathcal{F}^{\prime}=B(g,\tfrac{d}{2^{j-2}})\cap\mathcal{F}$ with cardinality of this set is bounded by $\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{d}{2^{j-2}},2c)$. Thus, we can apply Lemma 3.9 to (B.7), noting requiring $c>8$ implies $C>3$ as required. Moreover, the cardinality of $\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})$ is also bounded by $\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{d}{2^{j-2}},2c)$ by Lemma 3.2. Thus, continuing from (B.7), we have
@@
$\mathbb{P}(A_{j}\cap A_{j-1}^{c})$ $\leq\sum_{g\in\mathcal{L}(j-1)\cap B(\bar{f},\tfrac{d}{2^{j-2}})}3\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{d}{2^{j-2}},2c)\exp\left(-\Lambda(C,\sigma,L)\cdot n\delta^{2}\right)$
@@
$\leq 3\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{d}{2^{j-2}},2c)\right]^{2}\exp\left(-\Lambda(C,\sigma,L)\cdot n\delta^{2}\right)$
@@
$=3\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{j}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\right]^{2}\exp\left(-n\varepsilon_{j}^{2}\right).$
The last step used our definition of $\varepsilon_{j}$, $\delta$, and $C$.
Next, for $j=2$, note that $\mathbb{P}(A_{2}\cap A_{1}^{c})=\mathbb{P}(A_{2})$ since $\mathbb{P}(A_{1})=0$. Set $\delta=\tfrac{d}{2(C+1)}=\tfrac{d}{c}$. Well note that $\Upsilon_{2}$ is chosen from a maximal $\delta=\tfrac{d}{c}$-packing (hence covering) of $B(f_{R},d)\cap\mathcal{F}=B(\bar{f},d)\cap\mathcal{F}=\mathcal{F}$. Such a packing has cardinality bounded by $\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(d,c)\leq\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(d,2c)$, using Lemma 3.12. Again applying Lemma 3.9, we have
$\mathbb{P}(A_{2})$ $=\mathbb{P}(\|\Upsilon_{j}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1))$
$\leq 3\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(d,2c)\exp(-\Lambda(C,\sigma,L)\cdot n\delta^{2})$
$=3\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{2}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\exp(-n\varepsilon_{2}^{2})$
@@
$\leq 3\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{2}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\right]^{2}\exp(-n\varepsilon_{2}^{2}).$
Part 2: Deducing a bound on $\mathbb{P}(A_{J})$ for any $1\leq J\leq J^{\ast}$.
We now sum over our bounds using the $\mathbb{P}(A_{J})\leq\sum_{j=2}^{J}\mathbb{P}(A_{j}\cap A_{j-1}^{c})$ fact, noting that by Lemma 3.12,
$\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{j}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\leq\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{J}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c).$
Taking $a_{J}=\exp(-n\varepsilon_{J}^{2})$, we obtain
$\mathbb{P}(A_{J})$ $\leq 3\sum_{j=2}^{J}\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{j}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\right]^{2}\exp(-n\varepsilon_{j}^{2})$
$\leq 3\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{J}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\right]^{2}\sum_{j=2}^{J}\exp(-n\varepsilon_{j}^{2})$
@@
$\leq 3\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{J}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\right]^{2}\cdot\frac{a_{J}}{1-a_{J}}.$
Since $J^{\ast}>1$ and $1\leq J\leq J^{\ast}$, it follows (3.4) holds for $J$. Well, the first lower bound in (3.4) implies $\left[\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\tfrac{c\varepsilon_{J}}{\sqrt{\Lambda(\tfrac{c}{2}-1,\sigma,L)}},2c)\right]^{2}\leq\exp(n\varepsilon_{J}^{2}/2)$ and the second lower bound implies $\tfrac{1}{1-a_{J}}\leq 2$. Therefore,
$\mathbb{P}(A_{J})\leq 6\exp(n\varepsilon_{J}^{2}/2)\cdot a_{J}=6\exp(-n\varepsilon_{J}^{2}/2).$
For the following parts, let $B_{j}$ be the event $\{\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}>\kappa\varepsilon_{j}\}$ where $\kappa=\tfrac{1+3c}{2\sqrt{\Lambda(c/2-1,\sigma,L)}}$ and $f^{\ast}$ is the output of precisely $J^{\ast}-1$ steps of the algorithm, i.e., $\Upsilon_{J^{\ast}}$. Recall $f^{\ast\ast}$ is the output of $J^{\ast\ast}\geq J^{\ast}$ steps, i.e., $\Upsilon_{J^{\ast\ast}+1}$.
Part 3: Bounding $\mathbb{P}(B_{J})$ for integers $-\infty<J\leq J^{\ast}$.
First consider $1\leq J\leq J^{\ast}$. We can show $\mathbb{P}(B_{J})\leq\mathbb{P}(A_{J})$ for such $J$ since if $B_{J}$ occurs, using the triangle inequality and Lemma 3.3,
$\kappa\varepsilon_{J}$ $<\|\bar{f}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}+\|\Upsilon_{J}-f^{\ast}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}+\tfrac{d(2+4c)}{c2^{J}},$
which rearranging and substituting in $\kappa$ and $\varepsilon_{J}$ implies
$\|\bar{f}-\Upsilon_{J}\|_{L_{2}(\mathbb{P}_{X})}>\kappa\varepsilon_{J}-\tfrac{d(2+4c)}{c2^{J}}=\tfrac{d}{2^{J-1}}.$ (B.8)
Thus $B_{J}$ implies $A_{J}$, yielding
$\mathbb{P}(B_{J})\leq\mathbb{P}(A_{J})\leq 6\exp(-n\varepsilon_{J}^{2}/2).$
Noting then that $\kappa\varepsilon_{J}=\tfrac{(2+6c)d}{c2^{J}}>6d$ for $J\leq 0$, clearly $\mathbb{P}(B_{J})=0$ for all $J\leq 0$. Thus our bound in (B.8) remains valid all $-\infty<J\leq J^{\ast}$.
Part 4: Bounding $\mathbb{P}(\|\bar{f}-f^{\ast\ast}\|_{L_{2}(\mathbb{P}_{X})}>\kappa^{\prime}x)$ for $x\geq\varepsilon_{J^{\ast}}$ and $\kappa^{\prime}=(2+\tfrac{1+2c}{1+3c})\kappa$.